Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

Joshua and Jose

Expert replies
by gmatv09 » Mon Nov 09, 2009 10:55 pm
GMAT Prep Question...

OA A
You do not have the required permissions to view the files attached to this post.
Join the discussion
Source: — Problem Solving |

by life is a test » Tue Nov 10, 2009 7:06 am
6 people in total so each can be chosen with a probability of 1/6 -> prob of both of them being chosen is 1/6 * 1/5 --> the order in which they can be chosen can vary in two ways hence ans = 1/6 * 1/5 = 1/15

another way of looking at it is that the first person can be chosen in 2/6 ways and then second person can be chosen in 1/5 ways since the first one has already been picked -> again = 1/15

hope that helps.
Join the discussion

by gmatv09 » Tue Nov 10, 2009 2:28 pm
got it thanks..
Join the discussion

by heshamelaziry » Tue Nov 10, 2009 7:01 pm
life is a test wrote:6 people in total so each can be chosen with a probability of 1/6 -> prob of both of them being chosen is 1/6 * 1/5 --> the order in which they can be chosen can vary in two ways hence ans = 1/6 * 1/5 = 1/15

another way of looking at it is that the first person can be chosen in 2/6 ways and then second person can be chosen in 1/5 ways since the first one has already been picked -> again = 1/15

hope that helps.

Can this be answered in the form of Total desired outcomes/total outcomes ?????????????
Join the discussion

by satish.nagdev » Tue Nov 10, 2009 7:49 pm
2 / 6P2, as it mentions randomly
Wounded by GMAT but not dead
Join the discussion

by heshamelaziry » Tue Nov 10, 2009 8:44 pm
What is 6p2 ?
Join the discussion

by satish.nagdev » Tue Nov 10, 2009 9:22 pm
permutations, 6P2 so 6!/(6-2)!

**Edited, to correct mistake its 6P2 not 6P4
Wounded by GMAT but not dead
Join the discussion

by heshamelaziry » Tue Nov 10, 2009 9:36 pm
This give 15 not 1/15. also, shouldn't it be 6!/2!(6-2)! ?
Join the discussion

by heshamelaziry » Tue Nov 10, 2009 11:04 pm
Got it. Number of ways of chooisng Jose is 1. Number of ways of choosing Josh is 1. Number of ways choosing any 2 of the 6 = 6! / 2! (6-2)!

Probability = Desired outcomes = 1*1 / 6! / 2! (6-2)! = 1 / 15
Join the discussion

by crisro » Sat Jun 18, 2011 10:39 am
combination of 6 by 2 is 15
so we have 15 possible pairs and we need 1
Join the discussion

by GMATGuruNY » Sun Jun 19, 2011 1:51 am
Joshua and Jose work at an auto repair center with 4 other workers. For a survey on health care insurance, 2 of the 6 workers will be randomly chosen to be interviewed. What is the probability that Joshua and Jose will both be chosen?

a)1/15
b)1/12
c)1/9
d)1/6
e)1/3
P(first person selected is Joshua or Jose) = 2/6. (6 total people, among them both Joshua and Jose.)
P(next person selected is Joshua or Jose) = 1/5. (5 people left, among them either Joshua or Jose -- but not both -- since one was selected as the first person.)

Since we want these events to happen together, we multiply the fractions:
2/6 * 1/5 = 1/15.

The correct answer is A.
Private tutor exclusively for the GMAT and GRE, with over 20 years of experience.
Followed here and elsewhere by over 1900 test-takers.
I have worked with students based in the US, Australia, Taiwan, China, Tajikistan, Kuwait, Saudi Arabia -- a long list of countries.
My students have been admitted to HBS, CBS, Tuck, Yale, Stern, Fuqua -- a long list of top programs.

As a tutor, I don't simply teach you how I would approach problems.
I unlock the best way for YOU to solve problems.

For more information, please email me (Mitch Hunt) at [email protected].
Student Review #1
Student Review #2
Student Review #3
Join the discussion

by crisro » Sun Jun 19, 2011 7:56 am
GMATGuruNY wrote,

P(first person selected is Joshua or Jose) = 2/6. (6 total people, among them both Joshua and Jose.)
P(next person selected is Joshua or Jose) = 1/5. (5 people left, among them either Joshua or Jose -- but not both -- since one was selected as the first person.)

Since we want these events to happen together, we multiply the fractions:
2/6 * 1/5 = 1/15.

The correct answer is A.

but,
what if the two persons are selected together and not in two steps?
Join the discussion

by GMATGuruNY » Sun Jun 19, 2011 10:43 am
crisro wrote:GMATGuruNY wrote,

P(first person selected is Joshua or Jose) = 2/6. (6 total people, among them both Joshua and Jose.)
P(next person selected is Joshua or Jose) = 1/5. (5 people left, among them either Joshua or Jose -- but not both -- since one was selected as the first person.)

Since we want these events to happen together, we multiply the fractions:
2/6 * 1/5 = 1/15.

The correct answer is A.

but,
what if the two persons are selected together and not in two steps?
When we're asked to determine the probability that events will happen together, it makes no difference whether the events happen at the same time or one after the other; the probability is the same.

Another example:
A jar contains 3 red marbles and 2 blue marbles. If two marbles are selected at the same time, what is the probability that both marbles will be red?

To solve, we can determine the probability that the "first" marble is red and that the "second" marble also is red:
P(first marble is red) = 3/5.
P(second marble is red) = 2/4.

Since both events must happen together, we multiply:
3/5 * 2/4 = 3/10.
Private tutor exclusively for the GMAT and GRE, with over 20 years of experience.
Followed here and elsewhere by over 1900 test-takers.
I have worked with students based in the US, Australia, Taiwan, China, Tajikistan, Kuwait, Saudi Arabia -- a long list of countries.
My students have been admitted to HBS, CBS, Tuck, Yale, Stern, Fuqua -- a long list of top programs.

As a tutor, I don't simply teach you how I would approach problems.
I unlock the best way for YOU to solve problems.

For more information, please email me (Mitch Hunt) at [email protected].
Student Review #1
Student Review #2
Student Review #3
Join the discussion

by Thouraya » Sat Sep 17, 2011 12:49 am
@GMATGuruNY,

What if I want to solve the marbles problem with the approach 1-Probability None?

PS: i can't solve the Joshua/Jose problem with this approach because we have more than two choices?

Thank you so much
Join the discussion

by aplavakarthik » Sat Sep 17, 2011 2:17 am
GMATGuruNY wrote:
Joshua and Jose work at an auto repair center with 4 other workers. For a survey on health care insurance, 2 of the 6 workers will be randomly chosen to be interviewed. What is the probability that Joshua and Jose will both be chosen?

a)1/15
b)1/12
c)1/9
d)1/6
e)1/3
P(first person selected is Joshua or Jose) = 2/6. (6 total people, among them both Joshua and Jose.)
P(next person selected is Joshua or Jose) = 1/5. (5 people left, among them either Joshua or Jose -- but not both -- since one was selected as the first person.)

Since we want these events to happen together, we multiply the fractions:
2/6 * 1/5 = 1/15.

The correct answer is A.

Hi

can it be considered this way selecting 2 ppl from 6 is 6c2.

probability of selecting only these 2 is 1/6c2
Join the discussion